OpenStudy (anonymous):

calculate the wavelength, in nanometers, of the spectral line produced when an electron in a hydrogen atom undergoes the transition from the energy level n=7 to n=2

OpenStudy (anonymous):

Use the Rydberg formula \[\frac{1}{\lambda} = R_{E} \left( \frac{1}{n_{f}^2} - \frac {1}{n_{i}^2} \right)\] RE is the Rydberg constant If you have to derive that instead of just using a formula it involves a few more steps.